### v2

# Randomized methods for dynamical low-rank approximation

Abstract
We introduce novel dynamical low-rank methods for solving large-scale matrix differential equations, motivated by algorithms from randomized numerical linear algebra. In terms of performance (cost and accuracy), our methods overperform existing dynamical low-rank techniques. Several applications to stiff differential equations demonstrate the robustness, accuracy and low variance of the new methods, despite their inherent randomness. Allowing augmentation of the range and corange, the new methods have a good potential for preserving critical physical quantities such as the energy, mass and momentum. Numerical experiments on the Vlasov-Poisson equation are particularly encouraging.
The new methods comprise two essential steps: a range estimation step followed by a post-processing step. The range estimation is achieved through a novel dynamical rangefinder method. Subsequently, we propose two methods for post-processing, leading to two time-stepping methods: dynamical randomized singular value decomposition (DRSVD) and dynamical generalized Nyström (DGN). The new methods naturally extend to the rank-adaptive framework by estimating the error via Gaussian sampling.
1 Introduction
In this paper, we are interested in solving autonomous matrix differential equations of the general form
$\accentset{\mbox{\large.}}{A}(t)=F(A(t)),\quad A(0)=A_{0}\in\mathbb{R}^{m\times n},$ (1)
for some vector field $F:\mathbb{R}^{m\times n}\rightarrow\mathbb{R}^{m\times n}$ and time interval $[0,T]$. Examples of such differential equations include the discretization of partial differential equations in physics [EL18, CE22, CH22, Ein24], uncertainty quantification [BCSK17, AL24, FL18], quantum chemistry [SLC+24, CLS23] and also machine learning [SZK+22, SZCT23, SHNT23]; further applications will be provided in the next paragraphs. We refer to equation (1) as the full order model, and when the parameters $m$, $n$ are large, computing a dense solution to the problem becomes both computationally and memory intensive and is sometimes not feasible. To address this issue, reduced order modeling techniques have grown in popularity. A particularly successful one is the dynamical low-rank approximation [KL07, EL18], which uses low-rank compressions to efficiently store and evolve the solution over time.
On the other hand, the domain of randomized numerical linear algebra has emerged and is growing in popularity. A key achievement in the domain is the rangefinder algorithm, also called HMT, which comes with good theoretical guarantees [HMT11]. A direct application of the rangefinder is the randomized singular value decomposition (randomized SVD) algorithm which allows to approximate the truncated SVD of a (large) matrix $A\in\mathbb{R}^{m\times n}$ at a reduced cost. The randomized SVD algorithm scales well since it only requires a few matrix-vector multiplications, and such operations are particularly fast on GPUs. The generalized Nyström (GN) scheme [W+14, TYUC17, Nak20] is also a fast method for approximating the SVD of low-rank matrices, it is numerically stable and more accurate than the randomized SVD.
In this paper, we propose new time-stepping methods to approximate equation (1) with low-rank approximations. The new methods’ construction is inspired from the rangefinder, the randomized SVD and the generalized Nyström. At the moment, we are not aware of a direct formulation of randomized methods in the context time-dependent problems of the form (1), this paper is therefore an original attempt to link the domain of dynamical low-rank approximation and the domain of randomized numerical linear algebra. The new methods show a good potential for various applications in physics, uncertainty quantification, quantum chemistry, and machine learning. Before defining the new methods, we introduce notations and key results in the two domains.
1.1 Preliminary on randomized numerical linear algebra
The rangefinder algorithm
The idea of using a randomized subspace iteration for finding the range goes back to [RST10]. Quickly after, the powered randomized rangefinder algorithm was proposed and analyzed in [HMT11], and the algorithm is now simply called rangefinder or HMT. This algorithm solves the problem of approximating the range of a matrix $A\in\mathbb{R}^{m\times n}$ with a prescribed target rank $r$ and oversampling parameter $p$. More precisely, it aims to find a matrix $Q\in\mathbb{R}^{m\times(r+p)}$ with orthogonal columns such that
$\left\|A-QQ^{T}A\right\|_{2}=\min!$
The rangefinder algorithm consists of the following steps:
  1. 1.
    Draw a random Gaussian matrix $\Omega\in\mathbb{R}^{n\times(r+p)}$.
  2. 2.
    Compute the product: $B=A\Omega$.
  3. 3.
    Extract the range: $Q=\mathrm{orth}(B)$.
The algorithm scales particularly well since it only requires the application of $A$ to $r+p$ random vectors. The range is then extracted from the resulting tall and skinny $m\times(r+p)$ matrix $B$. The approximation can be improved by performing a few power iterations, which we summarize as follows:
$Q=\mathrm{orth}\left((AA^{T})^{q}A\Omega\right),$ (2)
and numerical stability is achieved with an orthogonalization step after each application of $A$. The detailed procedure is given in [HMT11, Algorithm 4.4], and we briefly state here the main convergence theorem from [HMT11, Corrolary 10.10].
Theorem 1.1 (Average spectral error of the rangefinder).
Let $A\in\mathbb{R}^{m\times n}$ with singular values $\sigma_{1}\geq\sigma_{2}\geq\ldots$. Choose a target rank $r\geq 2$ and an oversampling parameter $p\geq 2$, where $r+p\leq\min\{m,n\}$. Draw an $n\times(r+p)$ random Gaussian matrix $\Omega$ and construct $Q=\mathrm{orth}\left((AA^{T})^{q}A\Omega\right)$. Then, the expected approximation error verifies
$\mathbb{E}\left\|A-QQ^{T}A\right\|_{2}\leq\left[\left(1+\sqrt{\frac{r}{p-1}}\right)\sigma_{r+1}^{2q+1}+\frac{e\sqrt{r+p}}{p}\left(\sum_{j>r}\sigma_{j}^{2(2q+1)}\right)^{1/2}\right]^{1/(2q+1)}.$
In practice, the rangefinder is typically used with a relatively small oversampling parameter (typically $p\leq r$) and one or two power iterations are enough to get an average error close to the best error possible and also get a small variance. While random Gaussian matrices are generally a good choice, it appears that sub-sampled randomized trigonometric transforms (SRTTs) are slightly more accurate [MT20] and sparse maps allow faster matrix-vector multiplications.
The randomized SVD algorithm
A direct application of the rangefinder algorithm is the randomized SVD algorithm [HMT11] which we briefly summarize below.
  1. 1.
    Estimate the range by computing $Q=\mathrm{Rangefinder}(A,r,p,q)$.
  2. 2.
    Compute the reduced SVD of an $(r+p)\times n$ matrix: $\tilde{U}\Sigma V^{T}=\mathrm{svd}(Q^{T}A)$.
  3. 3.
    Update the basis and truncate the approximation: $\mathcal{T}_{r}(A)\approx\mathcal{T}_{r}(Q\tilde{U}\Sigma V^{T})$.
The randomized SVD scales particularly well since it only requires a few matrix-vector multiplications, and then costly operations are performed on small rectangular matrices. The truncation step is cost-free since the input is already given by its SVD, which can be truncated by keeping the $r$ first singular vectors/values. The analysis follows directly from Theorem 1.1, see also [HMT11, Theorem 10.5].
Since then, the method became standard for computing the SVD of large-scale matrices, and has also been extended to a variety of problems, including the principal component analysis [RST10, GGH+22], matrix completion [FYL18, WLC23], kernel matrix approximation [LBKL14], generalized SVD [SHvBW21], dynamic mode decomposition [Sch22] regularization of inverse problems [XZ13], image compression with GPU [JL14], tensor structures [BYDW18, AAAAM+21], and many more.
Generalized Nyström schemes
Generalized Nyström (GN) schemes are an alternative to the randomized SVD algorithm. Let us introduce first Nyström’s scheme, originally proposed in [Nys30] and more recently popularized in [WS00]. Let $A\in\mathbb{R}^{n\times n}$ be a square positive semi-definite (PSD) matrix, that is $A\succeq 0$. Then, a rank-$r$ approximation is given by
$A\approx AX(X^{T}AX)^{\dagger}(AX)^{T},$
where $X$ is an $n\times r$ random sketch matrix. The scheme became popular in machine learning for working with kernel matrices [WS00, SS02, FBCM04], and has been extensively analyzed in [Git11, GM13]. Interestingly, it has been observed in [HMT11, Mar19] and studied in [GM13] that computing
$A\approx AQ(Q^{T}AQ)^{\dagger}(AQ)^{T},\quad Q=\mathrm{Rangefinder}(AX)$
outperforms the randomized SVD in most situations. Later on, Nyström’s scheme has been generalized to rectangular matrices. Early algorithms were proposed in [CW09, W+14, TYUC17], and we refer to [Nak20] for a complete description and analysis of the scheme. For the purpose of the paper, we are interested in a particular variant of the generalized Nyström algorithm which uses the rangefinder algorithm and a second oversampling parameter $\ell\geq 0$, and which we define as follows:
$A\approx\tilde{A}=(AW)\ \mathcal{T}_{r}(Q^{T}AW)^{\dagger}\ (Q^{T}A),\qquad Q=\mathrm{Rangefinder}(A,r,p,q),\quad W=\mathrm{Rangefinder}(A^{T},r,p+\ell,q).$ (3)
As pointed out in [Nak20], the order in which operations are done is important for the numerical stability of the algorithm. Following our notation, the truncation to the target rank $r$ is applied first, and then the pseudoinverse of the truncated matrix is computed. Computing the pseudoinverse of an SVD is cheap since $(U\Sigma V^{T})^{\dagger}~{}=V\Sigma^{-1}U^{T}$. Overall, the cost of this version of generalized Nyström is roughly double the cost of a randomized SVD.
1.2 Dynamical low-rank approximation
The dynamical low-rank approximation (DLRA) was originally proposed in [KL07] and in [NL08] as a technique for evolving problems using low-rank factorizations to compress data. Let us now introduce important notations, briefly recall the fundamentals and then do a quick overview of the literature on the topic. In a separated paragraph below, we further discuss the differences between the new methods proposed in this paper and existing DLRA integrators.
Notations and theory
We consider the set of fixed rank matrices
$\mathcal{M}_{r}=\{M\in\mathbb{R}^{m\times n}\mid\mathrm{rank}(M)=r\},$
which is in fact a smooth manifold [LL03] embedded into $\mathbb{R}^{m\times n}$. The dynamical low-rank approximation is essentially the time-dependent variational principle (TDVP) applied to the vector field of the full order model and restricted to the above manifold of low-rank matrices; the DLRA evolves the projected dynamics
$\accentset{\mbox{\large.}}{Y}(t)=\mathcal{P}_{Y(t)}F(Y(t)),\quad Y(0)=Y_{0}\in\mathcal{M}_{r},$ (4)
where $\mathcal{P}_{Y}$ denotes the $\ell_{2}$-orthogonal projection onto the tangent space $\mathcal{T}_{Y}\mathcal{M}_{r}$, here defined at a certain low-rank matrix $Y~{}\in~{}\mathcal{M}_{r}$. The analysis of the accuracy of the DLRA relies on the following two assumptions.
Assumption 1 (One-sided Lipschitz).
There exists a constant $\ell\in\mathbb{R}$ such that for all $X,Y\in\mathbb{R}^{m\times n}$ we have
$\langle X-Y,\,F(X)-F(Y)\rangle\leq\ell\left\|X-Y\right\|_{F}^{2}.$
Assumption 2 (Closeness to the tangent bundle).
There exists a small constant $\varepsilon_{r}>0$ and a suitable neighborhood $\mathcal{N}$ of the solution to the full order model (1) such that for all $Y\in\mathcal{M}_{r}\cap\mathcal{N}$ we have
$\left\|F(Y)-\mathcal{P}_{Y}\left[F(Y)\right]\right\|\leq\varepsilon_{r}.$
Note that the one-sided Lipschitz constant can be negative, e.g. on diffusive problems. Under these two assumptions, the solution $Y(t)$ to the DLRA compared to the solution $A(t)$ to the full order model verify the error bound
[…]
and we refer to [KV19] for a proof. The DLRA model given in (4) is equivalent to a system of three coupled differential equations [KL07]. Solving the coupled system is numerically ill-conditioned in the presence of small singular values, and therefore robust techniques have been proposed in the literature.
A quick overview of numerical integrators
As above mentioned, solving the projected differential equation (4) can be numerically ill-conditioned in the presence of small singular values. Remember that we do expect to have small singular values since we are interested in problems with good low-rank approximations. It is therefore crucial to have methods that are numerically robust to the stiffness induced by small singular values. We now give a short overview of the literature on robust numerical methods and some of their applications.
The projector-splitting integrator [LO14] was the first integrator proposed, and its robust convergence was proved in [KLW16]. The projector-splitting integrator has a backward step and is therefore not suited for parabolic problems, then bases updates and Galerkin (BUG) methods were proposed in [CL22, CKL22, CKL24, CEKL24, Kus24]. Higher-order methods based on low-rank explicit Runge–Kutta schemes were proposed in [KV19, LCK24, NR25], and low-rank exponential Runge–Kutta methods for stiff differential equations are studied in [CV23]. Low-rank implicit schemes have also been studied in [NQE25, YEHS23, AC25, SV24, RV23], and a parallel-in-time low-rank integrator based on the parareal algorithm is discussed in [CGV23]. Dynamical CUR decomposition techniques have also been proposed in [DPN+23, NAB24, PN25], these methods have often less theoretical guarantees but have shown excellent results in practice and seem to be easier to implement and computationally less intensive than DLRA techniques, even though being closely related. Many of these methods have natural extensions to higher-dimensional problems with imposed tensor structures such as Tucker, tensor trains (or matrix product states) and tree tensor networks, see for example [KL10, LRSV13, LVW18, CL20, CL22, CLS23, GB24, EMP24, CKLS24].
The literature on applications of the dynamical low-rank approximation is now quite extensive, so we try to give a quick overview and more references can be found within the following papers. The DLRA has been successfully applied to various problems arising in quantum chemistry [SLC+24, CKLS24] and in physics: kinetics equations [EL18, Ein19, EOP20], radiation therapy [KS23], the Su–Olson problem [BEKK24], fission criticality [SEKM25] and also to stochastic differential equations [KNZ24, NT25, NT23]. Key properties of the numerical integrators in the context of physics are studied in [EJ21, EOS23, KEC23, EKS23]. Recent research also show applications in machine learning, such techniques are called dynamical low-rank training (DLRT) and are studied in [SZK+22, SZCT23, SL24, SZC+24, SYS25, KSW25].
1.3 Contributions and outline
As discussed in the previous section, the numerical issues due to the small singular values have been addressed by several numerical integrators. However, the low-rank integration of stiff differential equations is still challenging as they impose strong stepsize restrictions. Robust integrators based on exponential methods have been proposed in [OPW19, CV23], but they are limited to stiff differential equations with a Sylvester-like structure, and little is known for the low-rank integration of stiff differential equations with more general structures. The analysis of BUG-based methods [CL22, CKL24, Kus24] and projected Runge–Kutta methods [KV19] rely on the Lipschitz constant of the function of the vector field and have large hidden constants which are also observed numerically.
The idea of bringing randomization to the dynamical low-rank approximation has been recently studied in [LCK24] where they propose randomized low-rank Runge–Kutta methods. The key idea of these methods is to use randomization (with the generalized Nyström formula) within explicit Runge–Kutta schemes to limit the rank growth in the stages. By doing so, they avoid the computation of projections onto the tangent space, which makes the schemes robust to the presence of small singular values, and in some cases more robust than the projected Runge–Kutta methods proposed in [KV19]. However, the approach is currently limited to explicit Runge–Kutta schemes and, as a consequence, the convergence depends again on the Lipschitz constant of the vector field, leading to large hidden constants when applied to stiff differential equations. In contrast to the projector-splitting and BUG integrators, low-rank explicit Runge–Kutta methods do not present a decomposition into substeps, which reduces the potential for parallelization and requires to implement these methods from scratch.
Our main motivation is to develop BUG-like methods that can be applied to stiff differential equations without imposing strong stepsize restrictions. To achieve this, we introduce two new randomized time-stepping methods with a structure similar to existing BUG-based integrators. As a consequence, the new methods have a good potential for parallelization (in particular the dynamical generalized Nyström) and existing BUG implementations can be easily modified to implement the new methods. Additionally, we present novel rank-adaptive time-stepping methods where a Gaussian sampling is performed to estimate the errors at little extra cost. Our numerical experiments indicate that the new methods often outperform existing DLR integrators in terms of cost and accuracy, and demonstrate particular effectiveness in handling stiff differential equations characterized by large Lipschitz constants. With augmentation of the range and corange, the new methods also have a good potential for preserving critical physical quantities. A comprehensive convergence analysis of these methods remains an avenue for future research.
The paper is organized as follows. In Section 2, we propose a new method that we call dynamical rangefinder, which can be used to estimate the range (or corange) of the solution of a matrix differential equation with a prescribed rank, and we also propose an alternative rank-adaptive method using a prescribed tolerance instead. We provide numerical experiments on a toy problem to show the accuracy of the estimated range, and we compare the new method with the rangefinder directly applied to the reference solution. Next, with the help of the new dynamical rangefinder, we propose in Section 3 two new time-stepping methods that we call respectively dynamical randomized SVD and dynamical generalized Nyström. We show that, when the solution to the full order model is exactly low-rank and that the estimated range contains the range of the full solution, then the new methods are exact. In Section 4, we apply the two new time-stepping methods to various problems on which we compare the performance (cost and accuracy) with different kind of DLRA techniques.
2 Dynamical randomized rangefinder
Recall the full order problem (1) defined by the differential equation
$\accentset{\mbox{\large.}}{A}(t)=F(A(t)),\quad A(0)=A_{0}\in\mathbb{R}^{m\times n}.$
A common notation for the numerical flow computed by a numerical solver is $\phi^{F}_{t}(A_{0})$ and $\varphi^{F}_{t}(A_{0})$ for the exact flow (or exact solution). In this paper, we are not studying the error of numerical flows but only considering exact flows. Since the dependence on the initial value is not relevant for us, we use the explicit notation $A(t)$ for denoting the exact flow of the full order problem at time $t$, that is the exact solution to the full order model.
Let us fix a stepsize $h\in(0,T]$ and a target rank of approximation $r$, we now focus on the problem of finding the range which approximates best the exact solution $A(h)$. In other words, we seek a matrix $Q_{h}\in\mathbb{R}^{m\times r}$ such that
$\left\|A(h)-Q_{h}Q_{h}^{T}A(h)\right\|_{2}=\min!$
The notation ”$\min!$” is similar to the notation used in [FLLN24]. In this section, we start by discussing an idealized method where the exact solution to the full order model is available. Then, we propose a practical method to find such a matrix $Q_{h}$ without having to solve the full order model, and we propose a way to improve the approximation with dynamical power iterations. Finally, we discuss a method for automatically choosing the approximation rank for a prescribed tolerance using an error criterion with low probability of failure.
2.1 An idealized dynamical rangefinder method
As it is sometimes done in the literature on the dynamical low-rank approximation (e.g. in [LO14] and [CL22]), we start by studying the problem on the exact flow $A(t)$, and we consider the following steps:
  • Sketch a random Gaussian matrix $\Omega\in\mathbb{R}^{n\times(r+p)}$.
  • Solve the differential equation
    {.$&\accentset{\mbox{\large.}}{\mathcal{B}}(t)=F(A(t))\Omega,\quad t\in[0,h]\\ &\mathcal{B}(0)=A_{0}\Omega.\end{aligned}\right.$ (6)
  • Extract the basis at final time: $\mathcal{Q}_{h}=\mathrm{orth}(\mathcal{B}(h))$.
When the rank of the solution is over-estimated, the following lemma shows that the above procedure is exact.
Lemma 2.1 (Exact range property).
Fix a sketch matrix $\Omega\in\mathbb{R}^{n\times(r+p)}$ and assume that $\operatorname{rank}(A(h))\leq r+p$. Then, the matrix $Q_{h}$ obtained from the above procedure verifies $\mathrm{Range}(Q_{h})=\mathrm{Range}(A(h))$ and therefore
$\left\|A(h)-\mathcal{Q}_{h}\mathcal{Q}_{h}^{T}A(h)\right\|_{2}=0.$
Proof.
Suppose that $\mathrm{rank}(A(h))=R\leq r+p$, then we can decompose
$A(h)\Omega=U_{h}\Sigma_{h}V_{h}^{T}\Omega,$
where $U_{h}\in\mathbb{R}^{m\times R}$ and $V_{h}\in\mathbb{R}^{n\times R}$ have orthonormal columns. Since $\Omega\in\mathbb{R}^{n\times(r+p)}$ is a Gaussian matrix, the result of the product $W=V_{h}^{T}\Omega\in\mathbb{R}^{R\times(r+p)}$ is also a Gaussian matrix and has therefore full row-rank with probability one. By definition, we get
$\mathcal{B}(h)=A_{0}\Omega+\int_{0}^{h}F(A(t))\Omega\ \mathrm{dt}=A_{0}\Omega+\left[A(h)-A(0)\right]\Omega=A(h)\Omega=U_{h}\Sigma_{h}W,$
and we conclude by noticing that the following spaces are identical:
$\mathrm{Range}(Q_{h})=\mathrm{Range}(\mathcal{B}(h))=\mathrm{Range}(U_{h})=\mathrm{Range}(A(h)).$
As pointed out in the course of the above proof, the idealized procedure is, in fact, the rangefinder algorithm applied to the exact solution evaluated at time $h$, that is $A(h)$. It is also known that the projector-splitting integrator [LO14] and the BUG integrator [CL22] satisfy the exactness property on the exact flow.
2.2 A practical dynamical rangefinder with dynamical power iterations
In practice, the exact flow is unknown and equation (6) cannot be solved. We therefore need to approximate it without solving the full order problem. To do so, we propose the dynamical rangefinder algorithm:
  1. 1.
    Sketch a random Gaussian matrix $\Omega\in\mathbb{R}^{n\times(r+p)}$ and compute its pseudoinverse $\Omega^{\dagger}=(\Omega^{T}\Omega)^{-1}\Omega^{T}$.
  2. 2.
    (B-step) Solve the $m\times(r+p)$ differential equation
    $&\accentset{\mbox{\large.}}{B}(t)=F(B(t)\Omega^{\dagger})\Omega,\quad t\in[0,h]\\ &B(0)=A_{0}\Omega.\end{aligned}\right.$ (7)
  3. 3.
    Extract the basis at final time: $Q_{h}=\mathrm{orth}(B(h))$.
The approximation can further be improved by performing a few power iterations by approximating alternatively the range and corange of the matrix $A(h)$. One iteration of the dynamical power iteration is done as follows:
  1. 4.
    Solve the $n\times(r+p)$ differential equation
    $&\accentset{\mbox{\large.}}{C}(t)=F(Q_{h}C(t))^{T}Q_{h},\quad t\in[0,h]\\ &C(0)=A_{0}^{T}Q_{h}.\end{aligned}\right.$ (8)
  2. 5.
    Extract the basis at final time: $W_{h}=\mathrm{orth}(C(h))$.
  3. 6.
    Solve the $m\times(r+p)$ differential equation
    $&\accentset{\mbox{\large.}}{B}(t)=F(B(t)W_{h}^{T})W_{h},\quad t\in[0,h]\\ &B(0)=A_{0}W_{h}.\end{aligned}\right.$ (9)
  4. 7.
    Extract the basis at final time: $Q_{h}^{(1)}=\mathrm{orth}(B(h))$.
To perform $q$ dynamical power iterations, one has to repeat steps $4$ to $7$ starting from $Q_{h}^{(1)}$, then $Q_{h}^{(2)}$, etc. until obtaining $Q_{h}^{(q)}$. The dynamical rangefinder algorithm is summarized in Algorithm 1.
Algorithm 1 Dynamical rangefinder
1:Vector field $F:\mathbb{R}^{m\times n}\rightarrow\mathbb{R}^{m\times n}$, initial value $A_{0}\in\mathbb{R}^{m\times n}$, stepsize $h>0$, target rank $r>0$, oversampling parameter $p\geq 0$, number of power iterations $q\geq 0$.
2:
3:Sketch a Gaussian random matrix $\Omega\in\mathbb{R}^{n\times(r+p)}$.
4:Compute $\Omega^{\dagger}=(\Omega^{T}\Omega)^{-1}\Omega^{T}$. $\triangleright$ Solve a small linear system
5:Compute $B(h)$ by solving the $m\times(r+p)$ differential equation $\triangleright$ B-step
$\accentset{\mbox{\large.}}{B}(t)=F(B(t)\Omega^{\dagger})\Omega,\quad B(0)=A_{0}\Omega.$
6:Extract the basis: $Q=\mathrm{orth}(B(h))$. $\triangleright$ e.g. by pivoted QR
7:for $i=1$ to $q$ do $\triangleright$ Power iteration loop (optional)
8:     Compute $C(h)$ by solving the $n\times(r+p)$ differential equation
$\accentset{\mbox{\large.}}{C}(t)=F(QC(t)^{T})^{T}Q,\quad C(0)=A_{0}^{T}Q.$
9:     Extract the basis: $W=\mathrm{orth}(C(h))$.
10:     Compute $B(h)$ by solving the $m\times(r+p)$ differential equation
$\accentset{\mbox{\large.}}{B}(t)=F(B(t)W^{T})W,\quad B(0)=A_{0}W.$
11:     Extract the basis: $Q=\mathrm{orth}(B(h))$.
12:end for
13:Return $Q\in\mathbb{R}^{m\times(r+p)}$
2.3 Adaptive dynamical rangefinder
In most applications, the rank is not known a priori. Instead, one might prefer to prescribe an error tolerance and let the method find an appropriate rank such that the error $\left\|A(h)-Q_{h}Q_{h}^{T}A(h)\right\|_{2}$ will be smaller than the tolerance. The so-called adaptive randomized rangefinder [HMT11, Algorithm 4.2] relies on a lemma from [WLRT08]:
Lemma 2.2 (Approximate norm estimate via Gaussian sampling).
Let $M\in\mathbb{R}^{m\times n}$, choose an integer $K\geq 1$ and a real number $\alpha>1$. Draw an independent family $\{\omega^{(i)}:i=1,2,\ldots,K\}$ of standard Gaussian vectors. Then,
$\left\|M\right\|_{2}\leq\alpha\sqrt{\frac{2}{\pi}}\cdot\max_{i=1,\ldots,K}\left\|M\omega^{(i)}\right\|_{2}$
with probability at least $1-\alpha^{-K}$.
Taking $M=A(h)-QQ^{T}A(h)$ and $\alpha=10$ leads to the following estimate:
$\left\|A(h)-QQ^{T}A(h)\right\|_{2}\leq 10\sqrt{\frac{2}{\pi}}\cdot\max_{i=1,\ldots,K}\left\|A(h)\omega^{(i)}-QQ^{T}A(h)\omega^{(i)}\right\|_{2}$
with probability at least $1-10^{-K}$. Recall that the dynamical rangefinder makes the approximation $B(h)\approx A(h)\Omega$ and therefore
$A(h)\omega^{(i)}=A(h)\Omega_{[:,i]}\approx B(h)_{[:,i]}.$
In other words, $B(h)$ can be used as an accurate error estimate at little extra cost. We define the adaptive dynamical rangefinder algorithm as follows:
  1. 1.
    Set the block sampling size $K$ according to the prescribed tolerance and probability of failure.
  2. 2.
    Initial estimation of the range: $Q_{h}=\mathrm{DynamicalRangefinder}(F,A_{0},h,K,p=0,q=0)$.
  3. 3.
    Draw a new random Gaussian matrix $\Omega\in\mathbb{R}^{n\times K}$.
  4. 4.
    Solve the $m\times K$ differential equation
    $&\accentset{\mbox{\large.}}{B}(t)=F(B(t)\Omega^{\dagger})\Omega,\quad t\in[0,h]\\ &B(0)=A_{0}\Omega.\end{aligned}\right.$
  5. 5.
    Compute the error estimate $E=\max_{i=1,\ldots,K}\left\|B(h)_{[:,i]}-Q_{h}Q_{h}^{T}B(h)_{[:,i]}\right\|_{2}$.
  6. 6.
    If the error is too large, augment $Q_{h}$ with the new basis elements in $B(h)$ and repeat from step 3.
The above procedure is detailed in Algorithm 2.
Algorithm 2 Adaptive dynamical rangefinder
1:Vector field $F:\mathbb{R}^{m\times n}\rightarrow\mathbb{R}^{m\times n}$, initial value $A_{0}\in\mathbb{R}^{m\times n}$, stepsize $h>0$, tolerance $\tau>0$, max. failure probability $\beta>0$.
2:
3:Set $K=-\lceil\log(\beta)/\log(10)\rceil$ and $\varepsilon=\sqrt{\frac{\pi}{2}}\cdot\frac{\tau}{10}$. $\triangleright$ Block size and tolerance
4:Set $r=K$ and $p=0$.
5:Compute $Q_{h}=\mathrm{DynamicalRangefinder}(F,A_{0},h,r,p,q=0)$.
6:Set $E=\infty$ and $j=1$
7:while $E>\varepsilon$ do
8:     Sketch a random Gaussian matrix $\Omega\in\mathbb{R}^{n\times K}$ and compute $\Omega^{\dagger}=(\Omega^{T}\Omega)^{-1}\Omega^{T}$.
9:     Compute $B(h)$ by solving the $m\times K$ differential equation $\triangleright$ B-step
$\accentset{\mbox{\large.}}{B}(t)=F(B(t)\Omega^{\dagger})\Omega,\quad B(0)=A_{0}\Omega.$
10:     Update the error
$E=\max_{i=1,\ldots,K}\left\|B(h)_{[:,i]}-Q_{h}Q_{h}^{T}B(h)_{[:,i]}\right\|_{2}.$
11:     if $E>\varepsilon$ then $\triangleright$ Augment the basis
12:         Set $j=j+1$ and update the basis: $Q_{h}=\mathrm{orth}([Q_{h},B(h)])$triangleright e.g. by a QR update
13:     end if
14:end while
15:Return $Q_{h}\in\mathbb{R}^{m\times jK}$
2.4 Finding the range: a toy experiment
Let us consider the following differential equation:
$\accentset{\mbox{\large.}}{X}(t)=W_{1}X(t)+X(t)+X(t)W_{2}^{T},\quad X(0)=D,$ (10)
where $W_{1},W_{2}\in\mathbb{R}^{n\times n}$ are two random antisymmetric matrices, and $D\in\mathbb{R}^{n\times n}$ is a diagonal matrix such that $D_{ii}=2^{-i}$. The closed form solution is given by
$X(t)=e^{tW_{1}}\ e^{t}D\ e^{tW_{2}^{T}},$ (11)
and can therefore be used to compare the idealized and practical methods above defined. In the following, we set the parameters to $n=100$ and $h=0.1$. On one hand, we perform the idealized method by applying the rangefinder algorithm to the closed form solution $X(h)$ and we extract the matrix $Q$ with orthogonal columns. On the other hand, we perform the new dynamical rangefinder algorithm applied to equation (10) and we extract the matrix $Q_{h}$ with orthogonal columns. We emphasize that in the dynamical rangefinder algorithm, we never solve the full order model but only the reduced differential equation as defined in (7). The target rank is $r=5$ and the oversampling parameter $p$ varies from 00 to $12$. Due to the random nature of the algorithm, we perform $100$ simulations and show a box-plot showing the relative errors made by each method compared to the reference solution
$e_{1}=\frac{\left\|X(h)-QQ^{T}X(h)\right\|_{F}}{\left\|X(h)\right\|_{F}},\quad e_{2}=\frac{\left\|X(h)-Q_{h}Q_{h}^{T}X(h)\right\|_{F}}{\left\|X(h)\right\|_{F}}.$
The results are shown in Figure 1. As we can see, the two methods have a similar average accuracy, and are close to the rank $(r+p)$ approximation. In both cases, the variance and median error are improved when a power iteration is performed. When the oversampling parameter $p$ is large enough, the dynamical rangefinder accuracy is limited by the solver precision (here scipy RK45 with tolerance set to $10^{-12}$). This behavior is similar to the limit set by the machine precision for the rangefinder.
Figure 1: Application to the toy problem (10). The target rank of approximation is $r=5$ and the figure shows the relative error (at time $h=0.1$) made by the methods as we increase the oversampling parameter. The boxes show the median, first and third quartiles and extreme values of the $100$ simulations.
3 Time-stepping methods
Recall again that we are solving the full order model (1) given by the differential equation
$\accentset{\mbox{\large.}}{A}(t)=F(A(t)),\quad A(0)=A_{0}\in\mathbb{R}^{m\times n}.$
In the following, we will assume that the problem can be well-approximated by a low-rank approximation, and in particular we will start from an approximate of the initial value $A_{0}\approx Y_{0}=U_{0}S_{0}V_{0}^{T}\in\mathcal{M}_{r}$ where $U_{0}\in\mathbb{R}^{m\times r}$ and $V_{0}\in\mathbb{R}^{n\times r}$ have orthogonal columns, and $S_{0}\in\mathbb{R}^{r\times r}$ is invertible. We could take for example $Y_{0}=\mathcal{T}_{r}(A_{0})$, or an approximate of it. We fix a stepsize $h>0$ and a target rank of approximation $r$, the goal is to find factors such that $A(h)\approx Y_{1}=U_{1}S_{1}V_{1}^{T}\in\mathcal{M}_{r}$ where $U_{1}$ and $V_{1}$ have orthogonal columns. Building on the newly defined dynamical rangefinder algorithm, we propose two new time-stepping methods which evolve such a low-rank approximation and which do not involve the projected dynamics (4). The new time-stepping methods are defined by one step, which has to be repeated to get an approximate solution on the whole time interval. Ultimately, we discuss a simple modification of the two methods that allows to update and automatically choose the rank at each time step.
3.1 Dynamical randomized SVD
We start by proposing a new method inspired by the randomized SVD which consists of two steps: (1) a range estimation via the dynamical rangefinder and (2) a post-processing step done on the dynamics projected onto the estimated range. We start by describing an idealized version of the algorithm where we use the exact solution of the full order problem, then we define the practical variant of the algorithm, and numerical experiments are presented in Section 4.
An idealized dynamical randomized SVD
Let us start by motivating the method with an idealized procedure. Consider a target rank $r$ and an oversampling parameter $p$ and let $\mathcal{Q}_{h}\in\mathbb{R}^{m\times(r+p)}$ be the matrix with orthogonal columns obtained from the idealized dynamical rangefinder method described in Section 2.1. We propose the following idealized procedure:
  • Solve the differential equation
    $&\accentset{\mbox{\large.}}{\mathcal{C}}(t)=F(A(t))^{T}\mathcal{Q}_{h},\quad t\in[0,h]\\ &\mathcal{C}(0)=A_{0}^{T}\mathcal{Q}_{h}.\end{aligned}\right.$ (12)
  • Compute a reduced SVD: $\tilde{U}_{h}\Sigma_{h}V_{h}^{T}=\mathrm{svd}(\mathcal{C}(h)^{T})$.
  • Update the approximation: $A(h)\approx\mathcal{A}_{1}=\mathcal{Q}_{h}\tilde{U}_{h}\Sigma_{h}V_{h}^{T}$.
Additionally, one can truncate the computed solution to the target rank $r$, but is is not strictly necessary here. A direct consequence of the exact range property stated in Lemma 2.1 is the exactness of the above procedure, which we state in the lemma below.
Lemma 3.1 (Exactness property).
Choose a target rank $r$, an oversampling parameter $p$, and let $\mathcal{A}_{1}$ be the approximation obtained from the above idealized procedure. If $\mathrm{rank}(A(h))\leq r+p$, then $\mathcal{A}_{1}=A(h)$.
Proof.
By definition, the solution to the differential equation (12) verifies
$\mathcal{C}(h)=A_{0}^{T}\mathcal{Q}_{h}+\left[A(h)^{T}-A_{0}^{T}\right]\mathcal{Q}_{h}=A(h)^{T}Q_{h}.$
Then, we get by construction that $A_{1}=\mathcal{Q}_{h}\tilde{U}_{h}\Sigma_{h}V_{h}^{T}=\mathcal{Q}_{h}\mathcal{Q}_{h}^{T}A(h)$ and the conclusion follows with Lemma 2.1, since then $\mathrm{Range}(\mathcal{Q}_{h})=\mathrm{Range}(A(h))$. ∎
Solving the differential equation (12) requires access to the solution to the full order model, and is therefore not feasible in practice. Instead, we propose to approximate the solution by projecting the differential equation on an approximate range.
A practical dynamical randomized SVD
Because of the time discretization, approximating equation (12) on the estimated range at time $h$ might lead to a poor numerical approximation. We therefore need to augment the approximation space with the information from the initial value. Doing so increases the computational cost of the method but leads to better numerical results. Starting from the initial approximation $A_{0}\approx Y_{0}=U_{0}S_{0}V_{0}^{T}\in\mathcal{M}_{r}$, we define one step of the dynamical randomized SVD (DRSVD) as follows:
  1. 1.
    Estimate the range at time $h$ and augment the space:
    $Q_{h}=\mathrm{DynamicalRangefinder}(F,Y_{0},h,r,p,q)\quad\leadsto Q=\mathrm{orth}([U_{0},Q_{h}]).$ (13)
  2. 2.
    (C-step) Solve the $n\times(2r+p)$ differential equation
    $&\accentset{\mbox{\large.}}{C}(t)=F(QC(t)^{T})^{T}Q,\quad t\in[0,h]\\ &C(0)=Y_{0}^{T}Q.\end{aligned}\right.$ (14)
  3. 3.
    Compute a reduced SVD of the $(2r+p)\times n$ matrix: $\tilde{U}_{h}\Sigma_{h}V_{h}^{T}=\mathrm{svd}(C(h)^{T})$, and project the range: $U_{h}=Q\tilde{U}_{h}$.
  4. 4.
    Truncate the approximated solution to the target rank: $A(h)\approx Y_{1}^{\mathrm{DRSVD}}=\mathcal{T}_{r}(U_{h}\Sigma_{h}V_{h}^{T})$.
The truncation is cost-free since the input matrix is already an SVD factorization. Moreover, all matrices can be stored by their low-rank factorization, which reduces the algorithm memory footprint. The most expensive steps are to solve the two differential equations involved, but such steps are typical in dynamical low-rank integrators (see e.g. [LO14, CL22, EL18]). The above procedure also verifies the exactness property with an extra assumption on the estimated range.
Lemma 3.2 (Exactness property of DRSVD).
Assume that the solution to the full order model (1) at time $h$ verifies $\mathrm{rank}(A(h))\leq r+p$, and assume that $A_{0}=Y_{0}=U_{0}S_{0}V_{0}^{T}$. Let $Q$ be as defined in (13). If for all $t\in[0,h]$,
$\mathrm{Range}(A(t))\subset\mathrm{Range}(Q),$
then the dynamical randomized SVD algorithm is exact: $A(h)=Y_{1}^{\mathrm{DRSVD}}$.
Proof.
Start by defining $L(t)=A(t)^{T}Q$ and note that $A(t)^{T}~{}=~{}A(t)^{T}QQ^{T}~{}=~{}L(t)Q^{T}$. In particular, it verifies the differential equation
$\accentset{\mbox{\large.}}{L}(t)=F(A(t))^{T}Q=F(QL(t)^{T})^{T}Q,\quad L(0)=A(0)^{T}Q.$
We therefore have $C(0)=L(0)$ and $\accentset{\mbox{\large.}}{C}(t)=\accentset{\mbox{\large.}}{L}(t)$ for all $t\in[0,h]$, which implies that $C(t)=A(t)^{T}Q$ and the conclusion follows since $Y_{1}^{\mathrm{DRSVD}}=\mathcal{T}_{r}(QC(h)^{T})=\mathcal{T}_{r}(QQ^{T}A(h))=\mathcal{T}_{r}(A(h))=A(h)$. ∎
The assumption made in the above lemma cannot be checked without having access to the exact solution of the full order model. However, we can already say that the assumption in satisfied in the limit case $(r+p)=\min\{m,n\}$. In practice, there are two possibilities to increase the probability of satisfying the assumption and get (close to) the exactness property: (i) increase the oversampling parameter $p$ and (ii) reduce the stepsize $h$. The detailed procedure is summarized in Algorithm 3.
Algorithm 3 Dynamical randomized SVD (DRSVD)
1:Vector field $F:\mathbb{R}^{m\times n}\rightarrow\mathbb{R}^{m\times n}$, initial value $Y_{0}=U_{0}S_{0}V_{0}^{T}\in\mathcal{M}_{r}$, stepsize $h>0$, target rank $r>0$, oversampling $p\geq 0$, number of power iterations $q\geq 0$.
2:
3:Stage A: Dynamical range estimation
4:Compute $Q_{h}=\mathrm{DynamicalRangefinder}(F,Y_{0},h,r,p,q)$triangleright See Algorithm 1
5:Augment the basis: $Q=\mathrm{orth}([U_{0},Q_{h}])$triangleright With pivoted QR, or else
6:
7:Stage B: Dynamical post-processing
8:Compute $C(h)$ by solving the $n\times(2r+p)$ differential equation $\triangleright$ C-step
$\accentset{\mbox{\large.}}{C}(t)=F(QC(t)^{T})^{T}Q,\quad C(0)=Y_{0}^{T}Q.$
9:Compute the reduced SVD of the $n\times(2r+p)$ matrix $C(h)^{T}=\tilde{U}\Sigma_{1}V_{1}^{T}$ and set $U_{1}=Q\tilde{U}$.
10:Return The SVD truncated to the target rank $r$: $A(h)\approx Y_{1}^{\mathrm{DRSVD}}=\mathcal{T}_{r}(U_{1}\Sigma_{1}V_{1}^{T})$.
3.2 Dynamical generalized Nyström
Similarly to the dynamical randomized SVD presented above, we could give an idealized generalized Nyström algorithm based on the exact solution, which would also verify the exactness property stated in Lemma 3.1. The ideas being very similar, we skip the discussion of the idealized method and present directly the practical algorithm. Consider a target rank $r>0$, two oversampling parameters $p$, $\ell\geq 0$, and a low-rank approximate of the initial value $A_{0}\approx Y_{0}=U_{0}S_{0}V_{0}^{T}\in\mathcal{M}_{r}$. We define one step of the dynamical generalized Nyström (DGN) algorithm as follows:
  1. 1.
    In parallel, estimate the range and corange, and augment the bases:
    $Q_{h}=\mathrm{DynamicalRangefinder}(F,Y_{0},h,r,p,q)$ $\leadsto Q=\mathrm{orth}([U_{0},Q_{h}]),$ (15)
    $W_{h}=\mathrm{DynamicalRangefinder}(F^{T},Y_{0}^{T},h,r,p+\ell,q)$ $\leadsto W=\mathrm{orth}([V_{0},W_{h}]).$
  2. 2.
    (BCD-step) In parallel, solve the three reduced differential equations for $t\in[0,h]$:
    $&\accentset{\mbox{\large.}}{B}(t)=F(B(t)W^{T})W,\\ &B(0)=A_{0}W,\end{aligned}\right.\qquad\left\{\begin{aligned} &\accentset{\mbox{\large.}}{C}(t)=F(QC(t)^{T})^{T}Q,\\ &C(0)=A_{0}^{T}Q,\end{aligned}\right.\qquad\left\{\begin{aligned} &\accentset{\mbox{\large.}}{D}(t)=Q^{T}F(QD(t)W^{T})W,\\ &D(0)=Q^{T}A_{0}W.\end{aligned}\right.$ (16)
  3. 3.
    Truncate and assemble the solution:
    $A(h)\approx Y_{1}^{\mathrm{DGN}}=B(h)\ \mathcal{T}_{r}(D(h))^{\dagger}\ C(h)^{T}.$ (17)
As for the generalized Nyström algorithm given in (3), the order in which the solution is assembled is important for the numerical stability of the method. We first truncate and only then efficiently compute the pseudoinverse of the truncated matrix. Similarly to the dynamical randomized SVD, the dynamical generalized Nyström algorithm verifies the following exactness property when we assume that the augmented range and corange contain respectively the range and corange of the exact solution.
Lemma 3.3 (Exactness property of DGN).
Assume that the solution to the full order model (1) at time $h$ verifies $\mathrm{rank}(A(h))\leq r+p$, and assume that $A_{0}=Y_{0}=U_{0}S_{0}V_{0}^{T}$. Let $Q$ and $W$ be the matrices as defined in (15). If for all $t\in[0,h]$,
$\mathrm{Range}(A(t))\subset\mathrm{Range}(Q)\quad\text{and}\quad\mathrm{Range}(A(t)^{T})\subset\mathrm{Range}(W),$
then the dynamical generalized Nyström algorithm is exact: $A(h)=Y_{1}^{\mathrm{DGN}}$.
Proof.
The argument made in the proof of Lemma 3.2 holds verbatim on the three differential equations (16), so we have the three quantities:
$B(h)=A(h)W,\quad C(h)^{T}=QA(h)^{T},\quad D(h)=Q^{T}A(h)W.$
The conclusion follows from the definitions and noticing that
$Y_{1}^{\mathrm{DGN}}=B(h)\ \mathcal{T}_{r}(D(h))^{\dagger}\ C(h)^{T}=A(h)W(Q^{T}A(h)W)^{\dagger}QA(h)^{T}=A(h),$
thanks again to the range inclusions and the assumption on the rank of $A(h)$. ∎
The detailed algorithm is summarized in Algorithm 4, and we refer to Section 4 for applications.
Algorithm 4 Dynamical generalized Nyström (DGN)
1:Vector field $F:\mathbb{R}^{m\times n}\rightarrow\mathbb{R}^{m\times n}$, initial value $Y_{0}=U_{0}S_{0}V_{0}^{T}\in\mathcal{M}_{r}$, stepsize $h>0$, target rank $r>0$, oversampling parameters $p\geq 0$ and $\ell\geq 0$, number of power iterations $q\geq 0$.
2:
3:Stage A: Dynamical range and corange estimation
4:Compute (in parallel): $\triangleright$ See Algorithm 1.
$Q_{h}=\mathrm{DynamicalRangefinder}(F,Y_{0},h,r,p,q)$
$W_{h}=\mathrm{DynamicalRangefinder}(F^{T},Y_{0}^{T},h,r,p+\ell,q)$
5:(Optional) Augment the two bases: $Q=\mathrm{orth}([U_{0},Q_{h}]),W=\mathrm{orth}([V_{0},W_{h}])$.
6:
7:Stage B: Solve three small differential equations in parallel $\triangleright$ BCD-step
8:Compute $B(h)$ by solving the $m\times(2r+p+\ell)$ differential equation
$\accentset{\mbox{\large.}}{B}(t)=F(B(t)W^{T})W,\quad B(0)=Y_{0}W.$
9:Compute $C(h)$ by solving the $n\times(2r+p)$ differential equation
$\accentset{\mbox{\large.}}{C}(t)=F(QC(t)^{T})^{T}Q,\quad C(0)=Y_{0}^{T}Q.$
10:Compute $D(h)$ by solving the $(2r+p)\times(2r+p+\ell)$ differential equation
$\accentset{\mbox{\large.}}{D}(t)=Q^{T}F(QD(t)W^{T})W,\quad D(0)=Q^{T}Y_{0}W.$
11:Stage C: Truncate and assemble according to Nyström’s formula
12:Compute a reduced SVD of the $(2r+p)\times(2r+p+\ell)$ matrix: $\tilde{U}\Sigma\tilde{V}^{T}=\mathrm{svd}(D(h))$.
13:Truncate it to the target rank: $\mathcal{T}_{r}(D(h))=\tilde{U}_{r}\Sigma_{r}\tilde{V}^{T}_{r}$.
14:Compute two QR decompositions and assemble:
$U_{1}R_{1}=B(h)\tilde{V}_{r},\quad V_{1}R_{2}=C(h)\tilde{U}_{r}\qquad\leadsto S_{1}=R_{1}\Sigma_{r}^{-1}R_{2}^{T}.$
15:Return $A(h)\approx Y_{1}^{\mathrm{DGN}}=U_{1}S_{1}V_{1}^{T}$.
3.3 Rank-adaptive time-stepping methods
In the two new methods above presented, a fixed target rank is prescribed, which assumes that we have a guess on the numerical rank of the solution. In practice, we do not have always such a guess and we typically need to change the rank at each time step. Unlike the projector-splitting integrator, the augmented space together with the oversampling parameter $p$ already allow some freedom in the truncation step so that the rank can change from one step to the next. However, the rank growth is then limited to $2r+p$, which might not be enough to catch fast changing solutions. Instead, we propose a more systematic approach based on the adaptive dynamical rangefinder algorithm proposed in Section 2.3, which allows to choose automatically the rank.
Consider a prescribed error tolerance $\tau>0$ and a certain probability of failure $\beta>0$. Let us denote the tolerance-based truncated SVD by $\mathcal{T}_{\tau}$ such that
$\left\|X-\mathcal{T}_{\tau}(X)\right\|_{2}\leq\tau\quad\text{ for any matrix }X\in\mathbb{R}^{m\times n}\text{ and tolerance }\tau\in\mathbb{R},$
and which can be computed by truncating the singular values of the matrix $X$. Starting from the initial low-rank approximate $Y_{0}=\mathcal{T}_{\tau}(A_{0})=U_{0}S_{0}V_{0}^{T}$, we define one step of the adaptive dynamical randomized SVD (ADRSVD) algorithm as follows:
  1. 1.
    Estimate the range according to $\tau$ and $\beta$, and augment the basis:
    $Q_{h}=\mathrm{AdaptiveDynamicalRangefinder}(F,Y_{0},h,\tau,\beta)\quad\leadsto Q=\mathrm{orth}([U_{0},Q_{h}]).$
  2. 2.
    (C-step) Solve the reduced differential equation (14) to get $C(h)$.
  3. 3.
    Compute a reduced SVD: $\tilde{U}_{h}\Sigma_{h}V_{h}^{T}=\mathrm{svd}(C(h)^{T})$, and project the range: $U_{h}=Q\tilde{U}_{h}$.
  4. 4.
    Truncate the approximated solution to the prescribed tolerance: $A(h)\approx Y_{1}^{\mathrm{ADRSVD}}=\mathcal{T}_{\tau}(U_{h}\Sigma_{h}V_{h}^{T})$.
The definition of the adaptive dynamical generalized Nyström (ADGN) is similar:
  1. 1.
    In parallel, estimate the range and corange according to $\tau$ and $\beta$, and augment the bases:
    $Q_{h}=\mathrm{AdaptiveDynamicalRangefinder}(F,Y_{0},h,\tau,\beta)$ $\leadsto Q=\mathrm{orth}([U_{0},Q_{h}]),$ (18)
    $W_{h}=\mathrm{AdaptiveDynamicalRangefinder}(F^{T},Y_{0}^{T},h,\tau,\beta)$ $\leadsto W=\mathrm{orth}([V_{0},W_{h}]).$
  2. 2.
    (BCD-step) Solve the three reduced differential equations (16) to get $B(h)$, $C(h)$, and $D(h)$.
  3. 3.
    Truncate up to the prescribed tolerance and assemble the solution:
    $A(h)\approx Y_{1}^{\mathrm{ADGN}}=B(h)\ \mathcal{T}_{\tau}(D(h))^{\dagger}\ C(h)^{T}.$ (19)
The implementation of the two adaptive methods above defined is very similar to the implementation given in Algorithm 3 and Algorithm 4, and we refer to Section 4.2 for an application of these two adaptive methods on the Allen–Cahn equation.
4 Applications
We now apply the new time-stepping methods to partial differential equations and we compare the performance (cost and accuracy) with state-of-the-art DLRA techniques. The implementation is done in Python, and the code is fully available on GitHub [^1] The URL is: https://github.com/BenjaminCarrel/randomized-dynamical-low-rank for reproducibility. All experiments have been executed on a MacBook Pro with an M1 processor and 16GB of RAM.
4.1 A stiff differential equation with Lyapunov structure
Stiff differential equations are numerically challenging to solve as they impose a strong restriction on the stepsize. Methods like the projector-splitting integrator [LO14] and explicit (low-rank) Runge–Kutta methods [KV19, LCK24] are not suited for such problems. We consider here the heat equation with a constant source, given by the partial differential equation
$u(x,y;t)=\Delta u(x,y;t)+s(x,y),\quad u(x,y;0)=u_{0}(x,y),$ (20)
with Dirichlet boundary conditions and where $(x,y)\in\Omega$ and $t\in[0,T]$. Here, the source $s(x,y)$ is constant in time. A spatial discretization leads to the Lyapunov differential equation [UV20] defined by
$\accentset{\mbox{\large.}}{A}(t)=LA(t)+A(t)L+\alpha\frac{C}{\left\|C\right\|_{F}},\quad A(0)=A_{0},$ (21)
where $A(t)\in\mathbb{R}^{n\times n}$, $\alpha\geq 0$ and $t\in[0,T]$. For our numerical experiments, we use $\Omega=[-\pi,\pi]^{2}$ discretized with $n=256$ points. The discrete Laplacian operator is $L=(n-1)^{2}\cdot\mathrm{diag}(1,-2,1)\in\mathbb{R}^{n\times n}$. The construction of the initial value and the source is similar to [UV20, LCK24]. Since Assumption 2 is not satisfied at the initial value, we propagate the solution from time 00 to $10^{-4}$ and start our numerical experiments from there. We point out that Assumption 2 is only required for DLRA techniques, and we do not require it for the new methods. The new methods do not require this extra propagation of the initial value.
In Table 1, we compare the accuracy of the methods applied to equation (21) with a target rank $r=5$. When the stepsize is large (here $h=0.1$), existing DLRA methods perform poorly because the problem is stiff. The new methods perform well in comparison with a good accuracy and low variance despite the randomness. Note that the DGN algorithm with $q=1$ power iteration is able to recover the reference best rank approximation.
In Figure 2, we compare the time of computation of the new methods with the augmented BUG integrator. We can see here the large hidden constant of the augmented BUG integrator: with a stepsize of $h=0.1$ the method has not converged. We therefore need a smaller stepsize, here $h=0.01$, to get convergence of the augmented BUG. We point out that, for the same accuracy, the new methods are faster than the augmented BUG and the DGN can reach a higher accuracy when the rank of approximation is large enough.
(a) Existing DLRA methods with a rank of approximation $r=5$, and the reference solution truncated to the same rank.
Method Relative error after one step
Projector-splitting [LO14] (diverged)
Projected RK1 [KV19] 1.49e-01
Randomized low-rank RK1 [LCK24] 1.49e-01
Projected exponential RK1 [CV23] 1.01e-04
BUG integrator [CL22] 3.37e-05
Augmented BUG integrator [CKL22] 1.04e-06
Reference truncated to the target rank ($r=5$) 4.50e-09
Method $p=0$ $p=2$ $p=5$ $p=10$
DRSVD ($q=0$) 3.11e-04 1.93e-04 1.29e-04 8.29e-05
(Quartiles) [2.83e-04, 3.31e-04] [1.84e-04, 1.94e-04] [1.25e-04, 1.34e-04] [7.98e-05, 8.69e-05]
DRSVD ($q=1$) 3.25e-08 6.94e-09 6.08e-09 4.50e-09
(Quartiles) [3.25e-08, 3.26e-08] [6.41e-09, 7.16e-09] [4.71e-09, 7.22e-09] [4.50e-09, 4.50e-09]
DGN ($q=0$) 5.19e-09 4.66e-09 4.54e-09 4.51e-09
(Quartiles) [5.01e-09, 5.32e-09] [4.65e-09, 4.67e-09] [4.53e-09, 4.55e-09] [4.51e-09, 4.52e-09]
DGN ($q=1$) 4.50e-09 4.50e-09 4.50e-09 4.50e-09
(Quartiles) [4.50e-09, 4.50e-09] [4.50e-09, 4.50e-09] [4.50e-09, 4.50e-09] [4.50e-09, 4.50e-09]
(b) Median, first and third quartiles $([25\%,75\%])$ relative error ($30$ simulations) made by the new methods: dynamical randomized SVD (DRSVD) and dynamical generalized Nyström (DGN) with oversampling parameter $p$ and $q$ power iterations. For DGN, the second oversampling parameter is $\ell=0$.
Table 1: Relative error on the stiff matrix differential equation (21) after one step with $h=0.1$. The reference solver is a second-order exponential Runge–Kutta method with stepsize $\delta t=10^{-4}$, and the substeps were solved with the reference solver.
Figure 2: Accuracy and computation time (in seconds) of the methods applied to the Lyapunov equation (21). The integers next to the markers correspond to the target rank. The reference solver is a second-order exponential Runge–Kutta method with stepsize $\delta t=10^{-4}$, and the substeps were solved with the reference solver.
4.2 Reaction-diffusion with the Allen–Cahn equation
Another example of a stiff differential equation comes from the process of phase separation in multi-component alloy systems, described by the Allen–Cahn equation [AC72, AC73]. In its simplest form, it is given by
$\frac{\partial f}{\partial t}=\varepsilon\Delta f+f-f^{3},$ (22)
where $\Delta f$ is the diffusion term and $f-f^{3}$ is the reaction term. Dynamical low-rank techniques tailored for stiff differential equations have been successfully applied to this problem, see [RV23, CV23]. The stiffness can be controlled with the small parameter $\varepsilon$, and here we take $\varepsilon=0.01$ to get a moderately stiff problem. We consider a setting similar to [RV23] with initial value
$f_{0}(x,y)=\frac{2\ e^{-\tan^{2}(x)}\sin(x)\sin(y)}{1+e^{|\csc(-x/2)|}+e^{|\csc(-y/2)|}},$
the domain is $\Omega=[0,2\pi]^{2}$ and the spatial mesh contains $128\times 128$ points with periodic boundary conditions. After finite differences, we obtain the matrix differential equation
$\accentset{\mbox{\large.}}{X}(t)=AX(t)+X(t)A+X(t)-X(t)^{*3},$ (23)
where the power is taken element-wise. We consider the time interval $[0,T]=[0,10]$, and most of the reaction occurs between $t=3$ and $t=7$. At the end of the time interval, the solution has almost reached the steady-state. At any given time, the singular values of the solution decay exponentially fast.
The reaction component leads to a non-trivial evolution of the singular values, it is therefore a good problem for testing our new rank-adaptive methods. In Figure 3, we compare the new time-stepping methods with the rank-adaptive BUG integrator [CKL22]. We perform $30$ simulations of the random methods and show, at each time step, a boxplot showing the median and quartiles of the relative error. For a stepsize $h=0.5$, the error due to the time discretization is large for the adaptive BUG and adaptive DRSVD. Even though the variance of adaptive DRSVD is large, the method is always more accurate than the adaptive BUG on this example. On the other hand, the adaptive DGN has a tiny variance and a relative error close to the prescribed tolerance. Despite the differences in the relative error, all three methods estimate a rank of approximation close to the reference.
Figure 3: Rank-adaptive methods applied to the Allen–Cahn equation (23) with stepsize $h=0.25$ and $30$ simulations. The prescribed tolerance is $\tau=10^{-8}$. The reference is computed with a second order exponential Runge–Kutta method with time step $\delta t=0.005$, and the substeps in the low-rank methods are also solved with the reference solver.
4.3 Stochastic problems with the Burgers equation
In uncertainty quantification, the discretization of stochastic partial differential equations leads to large matrix differential equations. Let us consider here the one-dimensional Burgers equation
$\partial_{t}v+v\partial_{x}v=\nu\partial_{xx}v,\quad x\in\Omega,\quad t\in[0,T],$ (24)
with Dirichlet boundary conditions. As done in [NAB24], we sample the problem on several random initial conditions defined by
$v(x,0;\xi)=0.5\sin(2\pi x)\left(e^{\cos(2\pi x)}-1.5\right)+\sigma_{x}\sum_{i=1}^{d}\sqrt{\lambda_{x_{i}}}\psi_{i}(x)\xi_{i},\quad\xi_{i}\sim\mathcal{N}(0,1).$ (25)
More specifically, we draw $d$ random vectors $\xi_{i}\in\mathbb{R}^{s}$ that are normally distributed, and the eigenvalues $\lambda_{i}$ and eigenvectors $\psi_{i}(x)$ are extracted from the eigenvalue decomposition of the kernel matrix
$K_{ij}=\exp(-(x_{i}-x_{j})^{2}/2).$
The problem is discretized by associating each column to a random sample. Hence, taking the space $\Omega=[0,1]$ and using central finite differences leads to the matrix differential equation
$\accentset{\mbox{\large.}}{A}(t)=LA(t)-A(t)*(DA(t)),$ (26)
where the unknown matrix is $A(t)\in\mathbb{R}^{n\times s}$ and the operation * is the element-wise (Hadamard) product. The discrete differential operators are
$L=(n-1)^{2}\cdot\mathrm{diag}(1,-2,1)\in\mathbb{R}^{n\times n},\quad D=(n-1)/2\cdot\mathrm{diag}(-1,0,1)\in\mathbb{R}^{n\times n}.$
In our experiments, we set the parameters to $\nu=0.01,$ $\sigma_{x}=0.001$ and $d=4$. Then, we take $n=256$ spatial points, $s=64$ random samples, and solve the problem until $T=0.2$. For those parameters, the differential equation is not stiff. In Figure 4, we compare dynamical low-rank techniques on the stochastic Burgers equation (26) with a target rank of approximation $r=10$. Since the differential equation is not stiff, the projector-splitting works particularly well and has a small hidden constant compared to the BUG integrators. On this problem, the dynamical randomized SVD has a performance similar to the augmented BUG, and the dynamical generalized Nyström has the best accuracy with a relative error close to the best rank approximation.
Figure 4: Relative error over time of several methods applied to the stochastic Burgers equation (26) with stepsize $h=0.01$ and $30$ simulations. The reference solver is scipy RK45 method with absolute and relative tolerance set to $10^{-12}$, and the substeps are solved with the reference solver.
4.4 Kinetic equations with Vlasov-Poisson
The following Vlasov-Poisson equation [EL18] describes the evolution of electrons in a collisionless plasma, in the electrostatic regime with constant background ion density:
$\partial_{t}f(t,x,v)+v\cdot\nabla f(t,x,v)-E(f)(t,x)\cdot\nabla_{v}f(t,x,v)=0,$ (27)
[…]
where $f(t,x,v)$ is the particle-density function with $x\in\Omega_{x}\subset\mathbb{R}^{d}$ and $v\in\Omega_{v}\subset\mathbb{R}^{d}$. One aspect of equation (27) is its large memory footprint since it requires to store $O(n^{2d})$ floating point numbers, where $n$ is the number of grid points. To overcome the memory limitations, low-rank techniques have been introduced in [EL18] and we refer to references therein for state-of-the-art full rank solvers. The research on low-rank techniques for the Vlasov equation is still in development, see e.g. [GQ22, CE22, GQ24, UZ24]. Note that in the literature, the discretization is typically done after deriving equations for the low-rank factors of the function $f(t,x,v)$ itself. In this paper, we first perform a spatial discretization via finite differences and get the following matrix differential equation:
$\accentset{\mbox{\large.}}{A}(t)=-D_{x}A(t)V-E(t,A(t))A(t)D_{v},$ (28)
[…]
Linear Landau damping
The Landau damping is a phenomenon that occurs in plasma and can be modeled with equation (27). It is a damping of the electric field due to the interaction of the particles with the field. We consider the same setup as provided in [EL18], that is, the domain is $\Omega=(0,4\pi)\times(-6,6)$ with periodic boundary conditions and initial value
$f(0,x,v)=\tfrac{1}{\sqrt{2\pi}}\exp\left(-\tfrac{v^{2}}{2}\right)\left(1+\alpha\cos(kx)\right),$ (29)
where $\alpha=10^{-2}$ is the amplitude of the perturbation and $k=0.5$ is the wave number. For the discretization, we used $n_{x}=64$ and $n_{v}=256$ grid points in the $x$ and $v$ directions, respectively. For this specific problem, it is known that the theoretical rate of decay for the electric energy is $\gamma\approx-0.153$.
In Figure 5, we apply the new methods to the Landau damping problem until the final time $T=40$. The stepsize is $h=0.04$, and the substeps are solved with the reference solver. The projector-splitting (KSL2) is also given as a reference low-rank technique. The new methods accurately solve the problem and recover the theoretical rate until the tolerance of the reference solver is reached. At the end of the time interval, the modelling error induced by the new techniques is smaller than the modelling error induced by the projector-splitting integrator.
[…]
Two-stream instability
The two-stream instability [Bit13] is a well-known instability in plasma physics which is also modeled by equation (27). It is characterized by two counter-propagating beams of particles, which are perturbed by a small perturbation when $\alpha\neq 0$ in the initial value given below. The perturbation in the electric field grows exponentially, leading to the formation of structures in the plasma. Ultimately, the problem reaches the non-linear regime characterized by a nearly constant electric energy and filamentation of the phase space. While the linear regime admits good low-rank approximations, the non-linear regime is full rank and only energy, mass and momentum preservation can be expected from low-rank techniques. Again, we consider a setup similar to [EL18], the domain is $\Omega=(0,10\pi)\times(-6,6)$ and both directions are discretized with $n=128$ grid points and periodic boundary conditions. The initial value is
$f(0,x,v)=\frac{1}{2\sqrt{2\pi}}\left(e^{-(v-v_{0})^{2}}+e^{-(v+v_{0})^{2}}\right)\left(1+\alpha\cos(kx)\right),$ (30)
where $\alpha=10^{-3}$, $v_{0}=2.4$, and $k=1/5$.
We compare the new methods to the projector-splitting and augmented BUG integrators in Figure 6. In the linear regime, the new methods have errors comparable to the projector-splitting integrator, which is better than the augmented BUG integrator on this problem. In the non-linear regime, all methods keep the electric energy at a constant level. In Figure 7, we see that the new methods seem to approximate better the geometry of the reference solution even though the non-linear regime does not have a low-rank structure. The solution obtained by the DGN algorithm looks particularly symmetrical compared to other low-rank methods.
Figure 6: Numerical simulations of the two-stream instability (30) with stepsize $h=0.1$. The DRSVD is performed with oversampling parameter $p=5$ and $q=1$ power iteration. The DGN is performed with oversampling parameters $p=5$, $\ell=0$ and $q=1$ power iteration.
Figure 7: Two-stream instability solution at several time steps ($t=0,30,40,60$ from left to right). The solvers used are, from top to bottom: reference scipy RK45, projector-splitting, dynamical randomized SVD, dynamical generalized Nyström, augmented BUG. The color scale is shared for all plots.
Acknowledgements
Thanks to Bart Vandereycken and Daniel Kressner for their enthusiasm about this project, which encouraged me to pursue the research and publish the paper as single author. I also want to express my gratitude to Bart Vandereycken for his valuable feedback on the early drafts. This work was supported by the SNSF under research project 192363.